A uniform ladder AB of wieght W and length 2l rests with the end A against a smooth vertical wall and the end B on a rough horizontal ground. A man of weight equal to that of the ladder stands at the point C on the ladder, where BC = 5l/3 . The coefficient of friction between the ladder and the ground is 1/3. Given that the ladder is in limiting equilibrium when it makes an angle s to the horizontal;
(a) Show that tans = 2
(b) Find the magnitude of the reaction force at the wall and on the ground.
1
Expert's answer
2021-02-08T18:42:15-0500

"\text {let }O\text{ middle of ladder}"

"AO=OB= l;AC=\frac{l}{3};BC=\frac{5l}{3}"

"s-\text{angle s to the horizontal;}"

"F_{fr}-\text{friction force at point }B"

"T_A-\text{wall reaction force}"

"T_B-\text{reaction force at point B}"

"mg - \text{the gravity of the ladder at a point }O"

"mg - \text{the gravity of the man at a point }C"

"\text{forces along the X-axis (horizontal)}"

"T_A-F_{fr}=0"

"\text{forces along the Y-axis (vertical)}"

"T_B-mg-mg=0;T_B-2mg=0;"

"\text{equilibrium condition about point }B"

"T_A\sin{(s)}-mg*\cos(s)*AC-mg*\cos(s)*AO=0"

"T_A\sin{(s)}-mg*\cos(s)*\frac{1}{3}-mg*\cos(s)*l=0"

"T_A=\frac{4}{3}\cot(s)mg"

"F_{fr}=\frac{4}{3}\cot(s)mg"

"a)\text{friction static of the ladder and the man} - \frac{1}{3}mg+\frac{1}{3}mg=\frac{2}{3}mg"

"\text{ limiting equilibrium} F_{fr} \text{equal to static friction}"

"F_{fr}=\frac{4}{3}\cot(s)mg"

"\frac{4}{3}\cot(s)mg=\frac{2}{3}mg"

"\cot(s)=\frac{1}{2}"

"\tan(s)=2"

"b)" "T_A=F_{fr}=\frac{4}{3}\cot(s)mg"

"T_g-\text{reaction force on the ground.}"

"T_g= \sqrt{T_B^2+F_{fr}^2}"

"T_B=2mg"

"T_g= \sqrt{(4m^2g^2+\frac{16}{9}}\cot^2(s)m^2g^2)"


Answer:a)"\tan(s)=2" b)"T_A=\frac{4}{3}\cot(s)mg" ;"T_g= \sqrt{(4m^2g^2+\frac{16}{9}}\cot^2(s)m^2g^2)"